OpenStudy (anonymous):

I Need Quick Help.! Can Someone help? Solve for x: 3(x + 1) = -2(x - 1) - 4

OpenStudy (mertsj):

1. Distribute the 3 on the left 2. Distribute the -2 on the right 3. Collect like terms. 4. Add 2x to both sides 5. Subtract 3 from both sides 6. Divide both sides by 5

OpenStudy (anonymous):

what would it look like after i dsribute -2 on the right side?

OpenStudy (mertsj):

-2x+2-4

OpenStudy (anonymous):

i got 1.4 @Mertsj

OpenStudy (anonymous):

i think i did something wrong

OpenStudy (mertsj):

Post your work and I'll see if you did.

OpenStudy (anonymous):

3x+3 = -2x +2-4 3x+5 = -2x +3 5x + 5 = +2 -5 -5 5x/5 = 7/5

OpenStudy (mertsj):

3x+3 = -2x +2-4 What is 2-4?

OpenStudy (anonymous):

2

OpenStudy (mertsj):

2-4=-2 So on the right you have -2x-2

OpenStudy (mertsj):

And now the equation is 3x+3=-2x-2

OpenStudy (mertsj):

Now add 2x to both sides

OpenStudy (anonymous):

5x+3 = -2

OpenStudy (mertsj):

Subtract 3 from both sides

OpenStudy (anonymous):

5x= 1 ?

OpenStudy (mertsj):

Don't you have a calculator?

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):

ohh -5

OpenStudy (mertsj):

Now divide both sides by 5 Use your calculator

OpenStudy (anonymous):

-1

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (anonymous):

please /.\